A ball is thrown verically upwards with a speed of 60m/s.The distance it travels between t=0 s to t=8 s is? I know this question is extremely easy and I have even solved it.But the book is giving another answer. Please help.Its urgent

Grade:9

1 Answers

Manas Shukla
7 years ago
Applying the formula v = u + at , we can see the ball will reach its maximum height at t=6 secs if we take g = 10 m/s^2 . Then apply the formula v2= u2 + 2 a S to determine total displacement , whose value comes out to be 180 m.
Now for freefall of the ball we have 2 more seconds . So we apply S=ut + (½) at2 , whose answer is 20 m . Since total Distance is asked and not Displacement . We simply add 180+20 = 200m.